The Student Information Processing Board

The SIPB (pronounced �sip-bee�) is the Student Information Processing Board, a volunteer student group that centers around computing at MIT. A general description of SIPB’s job is to improve computing at MIT; which is something we try to accomplish in a number of ways.

SIPB provides a number of computer services to the MIT community. These range from helping users with walk-in or over-the-phone questions to maintaining a number of servers of use to the community, such as a web server (www.mit.edu), a Usenet news server (news.mit.edu), and an AFS cell (/afs/sipb.mit.edu). SIPB also maintains and develops Athena software for Linux and NetBSD, free operating systems that can be run on personal computers. Many MIT students choose to run one of these operating systems on their PC as it enables them to run almost all Athena software from home. (Athena is the computing environment that’s used in the public computer clusters at MIT)

SIPB is also an excellent place to learn more about computers. SIPB members tend to be most skilled with UNIX oriented operating systems and are always ready to impart their knowledge on those interested in learning. We have a library of useful books as well, available to anyone interested in taking a look.

The SIPB of�ce, in W20-557, is next to the Athena Cluster on the �fth �oor of the Student Center and is the place where most of SIPB’s activities are based. The of�ce is equipped with a number of different types of workstations: PC’s running Linux-Athena and NetBSD-Athena, Sun and SGI workstations like the ones in the public clusters, a few types of workstations which have long since been retired from the clusters, and a Mac which is currently the only publicly available Mac on campus. There is also a scanner available for the MIT community to use, as well as a Zip drive and a number of other drives that can be made available certain days of the week.

If you’re interested in the kinds of things the SIPB does, whether or not you know a lot about computers already, then you should consider becoming a SIPB prospective. A prospective is someone who is interested in becoming a SIPB member but has not yet been elected as a member. You can become a prospective simply by coming into the of�ce, �nding a SIPB member, and telling them you’re interested in becoming a prospective. Anyone with an interest in learning more about computers and also helping out the community is encouraged to become a prospective.

SIPB meetings are held every Monday at 7:30 pm in the SIPB of�ce, W20-557. Feel free to stop by then or any time if you’re interested in �nding out more about what we do. You can also reach us by calling x3-7788, or sending mail to sipb@mit.edu. We’re happy to answer your questions about SIPB or becoming a prospective, or to help you out with any computer-related questions you might have.